The angular displacement of a body is a function of time and is given by equation: tither= 10 + 3t+ 6t^2, where t is in seconds.
Determine the angular velocity, displacement and acceleration when t = 5 seconds.
1
Expert's answer
2020-08-26T11:19:31-0400

Angular velocity =

"\omega =\Delta\theta/\Delta t"

Differentiating "\Delta\theta" with respect to "\Delta t"

we get "\omega = 3 +12\times t"

when t=5

"\omega =" 3+12"\times t" = 3 +12"\times 5 = 63 rad/s^2"


displacement

"\theta=10 +3\times t +6\times t^2"

when "t=5"

"\theta=10 +(5\times 5)+(6\times 5^{2})"

"\theta=10+25+150 =175 rad"


"Acceleration""\alpha =\Delta\omega/\Delta t"

"Differentiating \Delta\omega" "/\Delta t"

we get "\alpha =12 rad/s^{2}"